### v1

# UniREditBench: A Unified Reasoning-based Image Editing Benchmark

Abstract
Recent advances in multi-modal generative models have driven substantial improvements in image editing. However, current generative models still struggle with handling diverse and complex image editing tasks that require implicit reasoning, underscoring the need for a comprehensive benchmark to systematically assess their performance across various reasoning scenarios. Existing benchmarks primarily focus on single-object attribute transformation in realistic scenarios, which, while effective, encounter two key challenges: (1) they largely overlook multi-object interactions as well as game-world scenarios that involve human-defined rules, which are common in real-life applications; (2) they only rely on textual references to evaluate the generated images, potentially leading to systematic misjudgments, especially in complex reasoning scenarios. To this end, this work proposes UniREditBench, a unified benchmark for reasoning-based image editing evaluation. It comprises 2,700 meticulously curated samples, covering both real- and game-world scenarios across 8 primary dimensions and 18 sub-dimensions. To improve evaluation reliability, we introduce multimodal dual-reference evaluation, providing both textual and ground-truth image references for each sample assessment. Furthermore, we design an automated multi-scenario data synthesis pipeline and construct UniREdit-Data-100K, a large-scale synthetic dataset with high-quality chain-of-thought (CoT) reasoning annotations. We fine-tune Bagel on this dataset and develop UniREdit-Bagel, demonstrating substantial improvements in both in-domain and out-of-distribution settings. Through thorough benchmarking of both open-source and closed-source image editing models, we reveal their strengths and weaknesses across various aspects.
Figure 1: UniREditBench covers both real-world and game-world reasoning scenarios across 8 primary dimensions and 18 sub-dimensions. We provide qualitative editing cases of (a) real-world multi-object interaction, and (b) game-world logical/strategy reasoning.
footnotetext: Equal contribution. Corresponding author.
1 Introduction
Table 1: Reasoning-based image editing benchmark comparison. Our UniREditBench excels in broader scenario and evaluation dimension coverage. “S-Obj” indicates single-object while “M-Obj” indicates multi-object.
Benchmark Size Reference Images Real World Scenario Game World Scenario Attribute Temporal Pose Spatial Motion Mechanic Medium Logical Long-planing Strategic Spatial (S-Obj) (S-Obj) (S-Obj) (M-Obj) (M-Obj) (M-Obj) (M-Obj) SmartEdit [11] 219 219 RISE [52] 360 70 KRIS [41] 1,267 50 UniREditBench 2,700 2,700
Recent advances in multimodal generative models have led to remarkable improvements in instruction-conditioned image editing. Generative models [40, 47, 50, 3, 33, 38, 44], including Step1X-Edit [20], FLUX-Kontext [2], Bagel [6], Nano Banana [7], and GPT-4o [12], have demonstrated a powerful ability to understand diverse textual instructions and generate semantically consistent image edits. In parallel, reinforcement learning-based training strategies [43, 18, 36, 31] are continuously advancing, further enhancing the capabilities of image editing models. With these rapid developments, the need for a more comprehensive benchmark to evaluate model editing capabilities across different aspects has become increasingly essential. Early benchmarks [20, 46] focus on local details or global stylistic changes, e.g., style transfer, color alteration, and object removal. However, they fail to cover editing tasks that require models to perform implicit reasoning [49, 8], which are commonly used in real-life applications. As illustrated in Fig. 1, when editing instructions involving real-world or human-defined game rules, current models often generate results that lack physical plausibility. To this end, recent efforts have introduced reasoning-aware evaluation across temporal, spatial, and logical dimensions [52], and proposed a knowledge-grounded taxonomy assessing factual, conceptual, and procedural knowledge types [41].
Despite their effectiveness, these benchmarks still face two significant challenges: (1) they primarily focus on single-object attribute changes in realistic scenarios, neglecting multi-object interactions and game-world scenarios involving human-defined rules (see Tab. 1). This narrow scope restricts their ability to evaluate how effectively models generalize across a wider range of complex reasoning contexts; Additionally, (2) they mainly rely on textual reference to evaluate the generated images [52, 41], which may lead to systematic misjudgments, especially in complex reasoning-based editing scenarios (see Fig. 2).
In this work, we posit that: (1) While current models exhibit proficiency in perceptual instruction following and simple reasoning editing settings (e.g., Transform an intact apple to a bitten one), they still struggle with complex reasoning-based image editing that necessitates the comprehension of multi-object interaction characteristics (e.g., Draw the paddle backward through the water) as well as logical constraints of puzzle and game scenarios (e.g., Control the player and push the box to the target), as illustrated in Fig. 1. (2) Relying solely on textual references in evaluating complex reasoning-based image editing task often leads to unreliable judgments. As shown in Fig.  2 (a), the text-reference-only evaluator assigns an inflated score even the edited image introduces an additional faulty path. Therefore, we intuitively believe that incorporating a ground-truth (GT) image as an additional visual reference can enable more precise evaluation.
Figure 2: Image editing evaluation comparison. Current text-reference-only evaluation potentially leads to misjudging, while our dual-reference evaluation results in more reliable assessments.
To this end, this work proposes UniREditBench, a unified benchmark for reasoning-based image editing assessment with broader evaluation dimension coverage and robust evaluation pipeline. Specifically, (1) we adopt a scenario-to-category hierarchical dimension design, covering diverse reasoning types in both real-world and game-world scenarios (shown in Fig. 1): it includes 2,700 carefully curated samples organized across 8 primary dimensions and 18 sub-categories, e.g., multi-object interaction in real world, and long-horizon game planning in game world. Meanwhile, (2) as illustrated in Fig. 2, in contrast to existing work that relies solely on textual references for evaluation, we introduce additional reference GT images to facilitate direct visual comparison with the generated image. By utilizing the visual cues provided by the reference image, the evaluator is able to able to more accurately and reliably assess the alignment of the generated image with the given instruction, as shown in Fig. 2 (b). Furthermore, to ensure the diversity and reliability of samples in this benchmark, we design a multi-scenario data synthesis pipeline. Specifically, as shown in Fig. 3, (a) For real-world scenarios, we first handcraft a few reference text prompts, including the original image description, the editing instruction, and the textual reference of edited effect. These prompts are then scaled up using the VLM. Finally, all resulted textual descriptions are directly used to generate pairs of original and edited image. (b) For game-world scenarios, we first design diverse game problems, and then use Python programs to generate image pairs, instructions, and textual reference of edited effects, ensuring both logical and visual correctness in these rule-intensive scenarios [16, 27]. Ultimately, all data samples in UniREditBench undergo VLM-based filtering and human inspection to ensure their reliability and accuracy.
Based on our data synthesis pipeline, we also propose UniREdit-Data-100K, a comprehensive reasoning-based image editing dataset with high-quality chain-of-thought (CoT) reasoning annotations, consisting of detailed, step-by-step reasoning traces generated using VLM, as shown in Fig. 3. To validate its reliability and effectiveness, we fine-tune the Bagel [6] on this dataset, resulting in UniREdit-Bagel. Experimental results demonstrate that the fine-tuned model achieves substantial improvements on both UniREditBench and other out-of-distribution benchmarks [52, 41]. Additionally, through comprehensive evaluation of both open- and closed-source editing models on our UniREditBench, we reveal their strengths and weaknesses across diverse reasoning-based scenarios.
Figure 3: Multi-Scenario data synthesis pipeline. (a) Real-world data synthesis pipeline; (b) Game-world data synthesis pipeline; and (c) Case study of our synthesized data.
Contribution: (1) We introduce UniREditBench, a unified benchmark for reasoning-based image editing that covers both real-world and game-world scenarios across 8 primary dimensions and 18 sub-dimensions, augmented with reference GT images to enable robust evaluation; (2) We design a multi-scenario data synthesis pipeline and develop UniREdit-Data-100K, a large-scale synthetic reasoning-based image editing dataset that includes high-quality CoT reasoning annotations. By fine-tuning the Bagel on this dataset, we develop UniREdit-Bagel and achieve substantial improvements, validating the effectiveness and reliability of our dataset; (3) Through comprehensive benchmarking of both open- and closed-source models, we systematically identify their strengths and weaknesses across diverse reasoning-based editing scenarios, offering valuable insights for advancing future models.
Figure 4: Qualitative cases of evaluation dimensions in UniREditBench. We present qualitative examples for each dimension across both real-world and game-world scenarios.
2 Related Work
Instruction-based Image Editing. Instruction-based image editing models aim to bridge semantic understanding of instructions with accurate visual manipulation. Traditional methods perform editing by altering the diffusion trajectory without requiring additional training, including partial denoising from intermediate SDE steps [21], cross-attention control [10, 30], mask-guided blending [38, 1, 33], CLIP- or diffusion-guided manipulation [14], and latent inversion for fidelity preservation [32, 13]. Besides, several studies employ visual-language models (VLMs) to provide prompts, spatial priors, or synthetic supervision to guide a generative editing model [3, 50, 51, 9]. Recent unified frameworks aim to use a single model for both image understanding and editing in a complementary direction [44, 40, 29]. For instance, Bagel [6] features a think mode that produces reasoning text prior to editing to enhance instruction fidelity and consistency. While effective, current methods still face challenges with complex reasoning-based editing, underscoring the need for comprehensive benchmarks to assess their performance across various reasoning scenarios.
Reasoning-based Benchmarks for Image Generation and Editing. In T2I generation, several benchmarks [34, 25, 17, 22] have been developed to assess the reasoning capabilities of models in generating images. For example, WISE [22] focuses on assessing models’ world knowledge, such as cultural and physical understanding, while UniGenBench++ [34] unifies semantic generation evaluation, covering 10 primary dimensions and 27 sub-dimensions, such as logic reasoning, relational understanding, supporting multilingual and varying-length assessments. In image editing evaluation, recent reasoning-based benchmarks like RISEBench [52] aim to examine temporal, spatial, and logical editing capabilities of editing models. Besides, KRISBench [41] introduces a knowledge-grounded taxonomy covering factual, conceptual, and procedural types. However, these benchmarks primarily focus on single-object knowledge and attribute reasoning. We suppose that extending evaluation to multi-object interactions and scenarios governed by human-defined rules is a crucial next step. As for image quality evaluation [45, 15], recent works like UnifiedReward [37, 35] adopt the “VLM-as-a-judge” paradigm, leveraging the powerful capabilities of VLMs to score and provide explanatory judgments. In image editing tasks, evaluation is more challenging because the evaluator needs to assess not only image quality but also understand complex editing instructions and final edited effects. Most studies like RISEBench and KRISBench utilize the property model [12], to rate instruction following, temporal consistency, and image quality. Despite effectiveness, their evaluation relies solely on textual references, which may lead to systematic misjudgments in complex reasoning tasks.
To this end, this work proposes UniREditBench, a unified reasoning-based image editing benchmark that spans a broad range of evaluation dimensions across real-world and game-world scenarios with multimodal dual-reference evaluation for more reliable and accurate assessments.
3 UniREditBench
3.1 Overview
With the rapid advancements in image editing models, existing benchmarks are gradually becoming less adequate to fully capture their comprehensive capabilities, particularly their reasoning-based editing abilities. Specifically, current benchmarks encounter two major challenges: (1) their evaluation primarily focuses on simple single-object attribute edits in real-world scenarios, neglecting complex multi-object interactions, as well as logical or strategic reasoning in game-world scenarios, where explicit human-defined rules govern the outcomes (Tab. 1); (2) their evaluation predominantly rely on clip-based metrics or VLM-based evaluators with text-only references, which may offer insufficient or inaccurate assessments, particularly in complex reasoning-intensive editing scenarios (Fig. 2).
To this end, this work proposes UniREditBench, a unified reasoning-based image editing benchmark that covers a broad spectrum of reasoning dimensions in different scenarios. Compared with previous studies, this benchmark exhibits several key superiorities:
  • Broader scenario and reasoning dimension coverage. It contains 2,700 high-quality samples organized into 8 primary reasoning dimensions and 18 sub-dimensions, spanning both real-world and game-world image editing tasks (Sec. 3.2).
  • Reliable dual-reference evaluation. For each sample assessment, we introduce both the textual reference and ground-truth (GT) image reference. This multi-modal reference enables vision-language model (VLM) evaluators to perform direct and fine-grained comparisons at both the textual and visual levels with the generated images, leading to more reliable evaluation (Sec. 3.3).
  • Scalable multi-scenario data synthesis. We propose an automatic data synthesis pipeline with distinct generation strategies tailored for real-world and game-world scenarios (Sec. 3.4).
3.2 Evaluation Dimensions
In real-life applications, image editing scenarios often involve diverse requirements spanning both real-world and game-world contexts, where complex contextual understanding and implicit reasoning capabilities are crucial for accurate image edits. Therefore, UniREditBench organizes reasoning-based image editing tasks into a scenario-to-category hierarchy framework. As illustrated in Fig. 1, it covers both real-world and game-world scenarios across 8 primary dimensions and 18 sub-categories, each representing a unique visual reasoning challenge with 150 human-inspected examples. We will elaborate on each dimension in the following.
3.2.1 Real-World Scenarios
Real-world scenarios involve editing tasks that reflect the perceptual and interaction dynamics commonly observed in natural environments. These tasks may involve transformations of individual objects or complex interactions among multiple objects. To handle such tasks, models must capture the semantic, physical, and temporal characteristics of objects, as well as their relationships.
  1. 1.
    Single-Object Transformation targets variations intrinsic to an individual object, including viewpoint and attribute changes that do not disrupt spatial relationships within the scene:
    • Viewpoint Transformation: Altering the perspective or viewing angle to exhibit alternative views of the same object (e.g., side, top-down, close-up).
    • Pose Adjustment: Modifying the articulation or positioning of an object’s parts, such as limb configurations or postural shifts.
    • Temporal Evolution: Simulating natural progressions over time like aging, decay, or seasonal changes impacting the object’s appearance.
    • Material Modification: Changing inherent surface or material properties (e.g., color, texture) while preserving geometry and location.
  2. 2.
    Multi-Object Interaction involves mutual influences and state changes arising from the physical or spatial interactions among multiple objects:
    • Structural Integrity Change: Physical deformations resulting from forces or collisions.
    • Motion State Change: Dynamics induced by contact or force transmission leading to altered movement or posture.
    • Mechanical Reaction: State transitions caused by device operation or functional interactions.
    • Medium Interaction: Changes mediated by substances or environmental factors that affect appearance or state.
    • Spatial Arrangement: Reorganization or repositioning of multiple objects within the scene.
Figure 5: Statistic visualization. We visualize (a) Word clouds and (b) data distribition of our UniREdit-data-100K.
3.2.2 Game-World Scenarios
Game-world scenarios consist of tasks within synthetic environments governed by human-defined rules, evaluating logical, strategic, spatial, and long-horizon reasoning capabilities. These tasks require models to plan, deduce, and act in accordance with the explicit rules that govern the environment.
  • Long-Horizon Planning requires multi-step sequential reasoning to accomplish distant goals, exemplified by navigation or puzzle games such as Maze-solving and Sokoban.
  • Logical Puzzle Solving involves constraint satisfaction and symbolic inference to produce valid solutions under formal rule sets, including Sudoku, Tic-Tac-Toe, and Word Search.
  • Strategic Reasoning requires resource management, adversarial planning over time, modeled after games like Pacman, Jewel2, and Space Invader.
  • Spatial Intelligence focuses on geometric and topological reasoning within 3D environments, such as reconstructing spatial layouts in gaming contexts.
Representative examples are provided in Figs. 1 and 4 to illustrate the scope and diversity of evaluation dimensions, and highlight the complexity and variety of tasks in our benchmark.
Figure 6: Qualitative editing result comparison. Our UniREdit-Bagel demonstrates significant superiority in both instruction following and visual quality compared with state-of-the-art closed-sourced and open-sourced models.
3.3 Dual-Reference Evaluation
Evaluating reasoning-based image editing is intrinsically challenging due to the need for the evaluator to accurately understand the implicit reasoning intentions within the instruction. To achieve reliable and comprehensive assessments, we introduce a VLM-based multi-dimensional scoring schema, leveraging both textual and visual evaluation references. Specifically, for each sample, this pipeline evaluates three core dimensions:
  • Instruction Following measures how accurately the generated image reflects the input instruction, focusing on whether the explicit effect of the edits is properly manifested. Here, the VLM compares the output image $G$ against both the textual reference of edited effect $R_{t}$ and the corresponding reference GT image $R_{i}$ to verify compliance:
    $S_{\mathrm{IF}}=\mathrm{VLM}(O,I,G,R_{i},R_{t})$
    where $O$ represents the original image, $I$ denotes the editing instruction.
  • Visual Consistency assesses the preservation of image regions and attributes unrelated to the edit instruction, ensuring that changes are localized and do not inadvertently alter irrelevant scene elements. This criterion favors models capable of accurate, fine-grained editing rather than wholesale regeneration:
    $S_{\mathrm{VC}}=\mathrm{VLM}(O,I,G)$
  • Visual Quality evaluates the realism and perceptual integrity of the generated output, checking for artifacts, distortions, and physical or logical implausibility in the final image:
    $S_{\mathrm{VQ}}=\mathrm{VLM}(G)$
    We choose GPT-4.1 [12] as VLM evaluator. Each score $S$ ranges from 1 to 5, following prior detailed scoring guidelines [52, 41]. Finally, the overall evaluation score aggregates these via weighted sum:
    $S_{\mathrm{Overall}}=\alpha_{1}S_{\mathrm{IF}}+\alpha_{2}S_{\mathrm{VC}}+\alpha_{3}S_{\mathrm{VQ}},$
    where $\alpha_{1}=0.5$, $\alpha_{2}=0.3$, and $\alpha_{3}=0.2$.
This setting prioritizes instruction following to emphasize the importance of accurately adhering to the instruction’s intent, and also incorporates visual consistency and quality, ensuring that areas unrelated to the instruction are preserved and that the overall image quality is maintained.
Table 2: In-domain quantitative comparisons on UniREditBench. GPT-4.1 is used as the evaluator. Best scores are in bold.
Model Real World Scenario Game World Scenario
Attribute Structure Physical Property Avg. Spatial Strategic Long-Horizon Logic Puzzle Avg. Overall
Modification Transform Interaction Response Intelligence Reason Plan Solving
[…]
FLUX-Kontext-Pro 45.68 45.65 42.85 40.17 43.59 43.54 44.03 49.80 40.07 44.36 43.72
Seedream4.0 69.54 73.13 67.88 62.40 68.24 39.27 43.54 43.79 51.91 44.63 57.05
Wan2.5 74.13 69.92 64.23 65.16 68.36 63.39 46.13 56.32 52.24 54.52 60.59
Nano Banana 77.54 78.45 72.41 77.39 76.45 66.26 56.83 56.35 65.85 61.32 68.38
GPT-4o 83.27 83.93 80.43 77.62 81.31 78.82 50.28 66.02 65.66 65.19 71.64
[…]
MagicBrush 43.10 46.24 43.88 47.70 45.23 63.43 33.77 30.83 35.17 40.80 40.98
Omnigen2 54.18 57.75 52.87 53.52 54.58 70.78 27.61 37.11 24.69 40.05 43.96
Step1X-Edit 59.50 57.37 56.07 56.75 57.42 62.90 34.17 44.63 44.03 46.43 50.12
Bagel-Think 61.27 59.83 53.95 57.11 58.04 65.65 43.46 43.82 40.18 48.28 51.25
Qwen-Image-Edit 75.19 73.16 71.80 67.54 71.92 57.03 34.80 47.37 37.70 44.22 56.46
UniREdit-Bagel (Ours) 79.24 78.61 76.43 71.81 76.52 84.93 73.98 86.14 84.01 82.27 78.87
3.4 Multi-Scenario Data Synthesis
Given the distinct characteristics of real- and game-world contexts, we develop specialized data generation process for each scenario, as illustrated in Fig. 3. Detailed elaboration of each data synthesis process is provided below.
For Real-World Scenario, we employ a “text-then-image” data generation strategy. Specifically, (1) this process begins with hand-crafted textual triples that describe the original image, the editing instruction, and the textual reference of edited effect (a reasoning-based narrative of the anticipated outcome). Next, we use the powerful VLM [5], to expand this initial set into a large corpus of text triples. Subsequently, (2) these curated textual triples are input to GPT-4o [12] to synthesize the original and edited images in alignment with the described textual reference of edited effect. (3) Finally, in the quality filtering stage, VLM [5] is used to assess the generated images based on visual fidelity, instruction alignment, and potential hallucination risks. Additionally, it generates reasoning chain-of-thought (CoT) text for each qualified instance, ensuring the production of high-quality, reasoning-based image editing training data.
In Game-World Scenario, game states are inherently well-suited to be represented as structured reasoning-based editing data, where instructions can naturally be solved using Python code. Inspired by Game-RL [48], (1) we first design a diverse collection of game-based problems and develop corresponding Python programs tailored to each category. (2) Then, these programs automatically generate paired original and edited images, along with instructions, textual reference effects, and programmatic CoT reasoning traces. (3) To bridge the gap between programmatic and natural language CoT reasoning formats, we use VLMs to convert these reasoning traces into explanations that align with human inference patterns. Finally, quality filtering are applied to ensure the integrity and reliability of the data.
Overall, this multi-scenario data synthesis pipeline generates our UniREditBench, a unified reasoning-based image editing benchmark, and UniREdit-Data-100K, a large-scale synthetic dataset with high-quality CoT annotations. We will provide a detailed elaboration of this dataset in the following section.
4 UniREdit-Data-100K
To enhance the capability of current generative models on reasoning-driven image editing, we propose UniREdit-Data-100K, which contains 100,421 samples spanning 8 reasoning dimensions and 18 categories defined in Sec. 3.2.
4.1 Statistical Analysis
UniREdit-Data-100K is designed with an emphasis on balance and diversity, ensuring that each reasoning category contains over 4,000 instances to effectively support model training across a wide range of editing tasks. It is divided into two primary scenario types: (i) Real-World Scenario, which captures natural object attributes and complex multi-object interactions, and (ii) Game-World Scenario, presenting structured, rule-based editing challenges, such as puzzles and strategic planning games. We visualize the word cloud for both real-world and game-world subsets in Fig. 5 (a) and the detailed distribution of samples across different categories in Fig. 5 (b). These visualizations highlight the extensive vocabulary of our dataset that captures the diverse visual attributes, as well as its broad coverage across various categories.
4.2 UniREdit-Bagel
To further validate the effectiveness of our dataset, we use it to fine-tune Bagel [6], a unified understanding and generative model. Specifically, each training sample consists of the input image $O$, an editing instruction $I$, a stepwise CoT text $C$ that grounds the edit effects step by step, and the target edited image $G$. During training, the original image and instruction are first input into the model, which then generates a textual reasoning trace and synthesizes the edited image. We supervise both the textual reasoning trace and the visual edit. Formally, for reasoning text supervision, we minimize the negative log-likelihood:
$\mathcal{L}_{\text{text}}=-\sum^{T}_{t}\log p_{\theta}\big(y_{t}\mid y_{<t},O,I\big).$
For image generation, we supervise the latent flow-matching loss [19] between the VAE latents of $O$ and $G$, conditioned on $(O,I,C)$:
$\mathcal{L}_{\text{img}}=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\;\big\|u_{\theta}(z_{t},t\,;\,O,I,C)-u^{\star}(z_{t},t)\big\|_{2}^{2},$
where $u_{\theta}$ is the learned time-conditioned velocity field on the latent path from $z_{O}$ to $z_{G}$, and $u^{\star}$ is the target velocity. Finally, the overall objective is
$\mathcal{L}=\lambda_{\text{text}}\mathcal{L}_{\text{text}}+\lambda_{\text{img}}\mathcal{L}_{\text{img}}.$
Under the influence of $\mathcal{L}_{\text{text}}$, the model enhances its reasoning ability through explicit CoT learning, which effectively guides the accurate image generation, while $\mathcal{L}_{\text{img}}$ improves both the correctness and fidelity of the edited image.
5 Experiment
5.1 Implementation Details
Baselines. We benchmark closed-source models including: GPT-4o [12], Nano Banana [7], Gemini-2.0 [5], Seedream4.0 [24], Wan 2.5 [28], and FLUX-Kontext-Pro [2], as well as open-source models including: Bagel [6], Qwen-Image-Edit [39], Step1X-Edit [20], FLUX.1-Kontext-dev [2], Emu2 [26], Omnigen2 [40], Omnigen [42], HiDream-Edit [4], MagicBrush [50], and AnyEdit [47].
Training and Evaluation. We train all components of Bagel [6] except the VAE model for 5,000 iterations using the Adam optimizer and the cosine learning rate scheduler on UniREdit-Data-100K. The scheduler includes 500 warm-up steps, with the peak learning rate of 2e-5 and minimum learning rate of 1e-6. During inference, we use the official inference settings provided by Bagel. To ensure fair comparisons with other baselines, we adopt the original inference configurations of these models.
5.2 Benchmarking Results on UniREditBench
As shown in Tab. 3.3, among closed-source models, GPT-4o achieves the highest average performance across all scenarios, with Nano Banana showing comparable capability. Wan2.5 delivers balanced results on real-world tasks but lags on game scenarios that require strategic reasoning. Besides, Seedream4.0 performs reasonably well on the structure transform dimension yet encounters challenges in game scenarios. Among open-source baselines, Qwen-Image-Edit performs strongly on real-world tasks such as attribute modification and structure transform. However, most models remain comparatively weak on game scenarios like Strategic Reasoning. Overall, compared with open-source methods, closed-source models, particularly GPT-4o, maintain a clear advantage. While some open-source models are competitive on specific real-world tasks, they generally struggle with complex reasoning in game scenarios. Notably, only GPT-4o and Nano Banana achieve an average score greater than 60 on game scenarios, underscoring that this setting remains highly challenging and serves as a useful test for current models.
5.3 Comparison Results of UniREdit-Bagel
Quantitative. UniREdit-Bagel achieves the best overall performance among all closed- and open-source models on UniREditBench, surpassing the second-place GPT-4o by a substantial margin. The largest gains occur in game-world scenarios (+17.08), indicating exceptional capability of understanding and processing complex reasoning image editing tasks. In out-of-distribution performance comparison, UniREdit-Bagel achieves the strongest open-source results across all four categories on RISEBench, shown in Tab. 5.3, improving upon the Bagel-Think baseline by 9.1 points and surpassing the closed-source Gemini-2.0-Flash-exp by 5.0 points. It also remains competitive with top closed-source models like Nano Banana and GPT-4o, narrowing the gap between open- and closed-source models.
Qualitative. The qualitative results presented in Fig. 6 highlight the strengths of our UniREdit-Bagel across various tasks. Specifically, in Fig. 6 (Row 4), most models fail to reliably reproduce the physical heat effect. Although several baselines, such as Nano Banana and Qwen-Image-Edit, successfully capture the heat-induced warping of a plastic bottle under sustained heat gun exposure, they fail to preserve the heat trace. Notably, UniREdit-Bagel not only renders the deformation accurately but also preserves the heat trace, offering superior visual consistency. Besides, in the Sokoban and Maze game settings (rows 1 and 3), Seedream-4.0, Nano Banana, and Wan-2.5 generally preserve instruction-irrelevant content but struggle with instruction-specific objectives. In contrast, UniREdit-Bagel excels in both fulfilling the instruction and maintaining the coherence of unrelated content.
Table 3: Out-of-distribution quantitative performance comparison on RISEBench [52]. GPT-4.1 is used as the evaluator. Best scores are in bold.
Models Temporal Causal Spatial Logical Overall
[…]
Gemini-2.0-Flash-pre 10.6% 13.3% 11% 2.3% 9.4%
Seedream-4.0 12.9% 12.2% 11.0% 7.1% 10.8%
Gemini-2.0-Flash-exp 8.2% 15.5% 23.0% 4.7% 13.3%
GPT-4o 34.1% 32.2% 37.0% 10.6% 28.9%
Nano Banana 25.9% 47.8% 37.0% 18.8% 32.8%
[…]
HiDream-Edit 0.0% 0.0% 0.0% 0.0% 0.0%
OmniGen 1.2% 1.0% 0.0% 1.2% 0.8%
Step1X-Edit 0.0% 2.2% 2% 3.5% 1.9%
Bagel 3.5% 4.4% 9.0% 5.9% 5.8%
FLUX.1-Kontext-Dev 2.3% 5.5% 13.0% 1.2% 5.8%
Qwen-Image-Edit 4.7% 10.0% 17.0% 2.4% 8.9%
Bagel-Think 4.7% 15.5% 14.0% 1.2% 9.2%
UniREdit-Bagel (Ours) 22.4% 18.9% 21.0% 10.6% 18.3%
6 Conclusion
This paper introduces UniREditBench, a unified reasoning-based benchmark for image editing with broader evaluation dimension coverage and a reliable dual-reference evaluation pipeline. Additionally, we design a multi-scenario data synthesis pipeline and release UniREdit-Data-100K, a large-scale dataset with high-quality chain-of-thought (CoT) annotations. To demonstrate its effectiveness, we fine-tune Bagel on this dataset, resulting in UniREdit-Bagel, which achieves significant improvements both quantitatively and qualitatively. Through comprehensive benchmarking of both open-source and closed-source image editing models, we highlight their strengths and weaknesses across various aspects.
Supplementary Material
Appendix A Data Filtering
We design a comprehensive, multi-stage pipeline that performs data filtering, i.e., instruction de-duplication, quality filtering, and human inspection, to remove redundancy and low-quality data. Detailed elaboration of the filtering pipeline is provided below.
A.1 Instruction De-duplication
During the first stage of the real-world scenario, text prompt for image editing are sampled from the Gemini-2.5-Pro, which may potentially introduce repeated or near-duplicate entries. We remove redundancy along two aspects: exact matches and semantic similarity.
  • Exact-Match Deduplication: We first normalize the original image description by converting it to lowercase and removing punctuation. Afterward, we extract the set of words from the normalized text. If two samples contain identical word sets, they are considered duplicates, as the descriptions are effectively the same. These duplicate samples are then filtered out to ensure data diversity.
  • Semantic-Similarity Deduplication: We use a sentence-transformers [23] model to extract sentence embeddings for both the original image description and edit instruction. We then compute the pairwise similarity between these embeddings. If the similarity score exceeds a threshold of 0.7 for either description and instruction, the samples are deemed semantically redundant and are filtered out to enhance dataset diversity.
These complementary exact-match and semantic filters improve dataset diversity by eliminating both literal and paraphrastic duplicates.
A.2 Quality Filtering
To ensure the quality of both the generated text and images, we evaluate and filter them across six key dimensions: text hallucination, instruction adherence, content preservation, visual quality, image hallucination, and CoT quality. Scores for each dimension are assigned by the Gemini-2.5-Pro on a 1–5 scale. Only samples that achieve the maximum score across all dimensions are retained.
  • Text Hallucination: We evaluate the textual reference for hallucinated content, defined as entities or visual effects that are not mentioned in the Instruction or that cannot be plausibly induced by the given Instruction.
  • Instruction Following: We compare the edited image with the textual reference to assess whether the generated visual changes accurately reflect the specified effects. Samples that demonstrate poor adherence to the instructions and text reference are discarded.
  • Content Preservation: We assess whether regions unrelated to the edit instruction, such as the background, remain consistent between the original and edited images, ensuring stability in unaffected areas.
  • Visual Quality: We assess whether the generated images meet fundamental quality standards, specifically by ensuring they are free from artifacts or degradation.
  • Image Hallucination: We examine the edited images for any unintended additions or alterations beyond the specified textual reference, such as the appearance of additional objects.
  • CoT Quality: We evaluate the correctness of the chain-of-thought (CoT) reasoning text, focusing on whether the analysis of the original image and instruction is logical and sound.
Figure 7: Benchmarking result visualization. (a) Closed-source model comparison; (b) Open-source model comparison.
Table 4: Quantitative comparisons on KRISBench [41]. GPT-4.1 is used as the evaluator. Best scores are in bold while second-best is underlined.
Model Attribute Spatial Temporal Social Natural Logical Instruction Overall
Perception Perception Prediction Science Science Reasoning Decompose Score
[…]
Doubao 70.92 59.17 40.58 65.50 61.19 47.75 60.58 60.70
[…]
Gemini-2.0 66.33 63.33 63.92 68.19 56.94 54.13 71.67 62.41
GPT-4o 83.17 79.08 68.25 85.50 80.06 71.56 85.08 80.09
[…]
MagicBrush 53.92 39.58 - 42.94 38.06 30.00 23.08 37.15
AnyEdit 47.67 45.17 - 38.56 42.94 36.56 26.92 38.55
Emu2 51.50 48.83 22.17 34.69 38.44 24.81 45.00 39.70
Step1X-Edit 55.50 51.75 - 44.69 49.06 40.88 22.75 43.29
Bagel-Think 69.27 67.58 - 65.00 62.11 47.33 49.22 60.77
UniREdit-Bagel (Ours) 71.75 71.00 - 69.20 65.99 59.91 51.55 65.45
Table 5: Detailed in-domain quantitative comparisons on UniREditBench. GPT-4.1 is used as the evaluator. Best scores are in bold.
Model Real World Scenario Game World Scenario Overall
Attribute Modification Structure Transform Physical Interaction Property Response Spatial Intelligence Strategic Reason Logic Puzzle Solving Long-Horizon Plan
Viewpoint Transformation Material Modification Pose Adjustment Temporal Evolution Structural Integrity Change Motion State Change Spatial Arrangement Mechanical Reaction Medium Interaction 3D Reconstruction Space Invader Jewel2 Pacman Word Search Tictactoe Sudoku Maze Sokoban
[…]
FLUX-Kontext-Pro 34.75 56.62 54.95 36.36 42.15 42.52 43.89 38.01 42.33 43.54 47.45 34.09 50.55 34.88 58.44 26.88 46.70 52.90 43.72
Seedream4.0 64.77 74.32 80.19 66.07 59.48 64.68 79.48 61.58 63.21 39.27 46.28 43.00 41.34 48.55 38.43 68.75 54.15 33.43 57.05
Wan2.5 72.97 75.30 79.97 59.86 64.51 66.56 61.60 63.67 66.66 63.39 44.67 53.73 40.00 58.54 60.59 45.94 65.17 47.47 60.59
Nano Banana 75.37 79.72 85.55 71.35 70.65 73.36 73.22 79.44 75.34 66.26 61.43 54.65 54.40 64.65 40.92 91.99 62.65 50.05 68.38
GPT-4o 83.55 82.98 92.16 75.70 76.10 76.32 88.88 78.97 76.28 78.82 58.88 44.95 47.02 64.55 63.80 68.62 82.27 49.77 71.64
[…]
MagicBrush 35.65 50.55 48.34 44.13 45.07 47.83 38.73 47.88 47.52 63.43 44.74 26.26 30.30 34.12 31.05 40.33 32.65 29.00 40.98
Omnigen2 50.32 58.03 67.73 47.77 49.90 57.63 51.09 55.05 51.98 70.78 51.17 1.50 30.17 23.07 46.17 4.85 39.63 34.58 43.96
Step-1X-Edit 51.82 67.18 63.29 51.45 54.98 62.18 51.03 55.72 57.78 62.90 33.62 35.40 33.48 43.02 49.92 39.17 54.53 34.73 50.12
Bagel-Think 58.38 64.15 63.38 56.27 54.40 58.63 48.82 56.47 57.75 65.65 47.30 42.12 40.97 47.80 40.35 32.40 48.83 38.80 51.25
Qwen-Image-Edit 72.08 78.31 81.75 64.57 68.78 69.51 77.12 67.83 67.25 57.03 36.58 37.79 30.02 48.47 33.00 31.63 60.02 34.71 56.46
UniREdit-Bagel (Ours) 84.43 74.06 85.13 72.08 74.31 71.95 83.05 71.17 72.45 84.93 86.93 61.25 73.76 87.48 70.65 93.90 97.73 74.55 78.87
A.3 Human Inspection
In addition to our automated filtering pipeline, we perform a final manual check of each data instance. To facilitate this, we developed two web-based interfaces and enlisted eight expert annotators to carry out two-stage filtering process:
  • Initial Filtering: Annotators remove samples with extremely erroneous textual references or substandard generated images.
  • Manual Correction: Annotators make refinements to the textual reference effect that are only slightly incorrect, ensuring alignment and accuracy.
Two web interfaces are shown in Figs. 11 and  12.
Appendix B Detailed Benchmarking Results
We provide detailed benchmarking results on our UniREditBench for each category in Tab. A.2.
Appendix C More Quantitative Results
We provide more quantitative out-of-distribution performance comparisons on KRISBench in Tab. A.2.
Appendix D More Qualitative Comparison Results
We provide additional qualitative comparisons on UniREditBench in Fig. 8 and 9, and comparisons on RISEBench in Fig. 10.
Figure 8: Qualitative editing result comparison on UniREditBench. Our UniREdit-Bagel demonstrates significant superiority in both instruction following and visual quality compared with state-of-the-art closed-sourced and open-sourced models.
Figure 9: Qualitative editing result comparison on UniREditBench. Our UniREdit-Bagel demonstrates significant superiority in both instruction following and visual quality compared with state-of-the-art closed-sourced and open-sourced models.
Figure 10: Qualitative editing result comparison on RISEBench. Our UniREdit-Bagel demonstrates significant superiority in both instruction following and visual quality compared with state-of-the-art closed-sourced and open-sourced models.
Appendix E Ethical statement
In this work, we affirm our commitment to ethical research practices and responsible innovation. To the best of our knowledge, this study does not involve any data, methodologies, or applications that raise ethical concerns. All experiments and analyses were conducted in compliance with established ethical guidelines, ensuring the integrity and transparency of our research process.
Figure 11: Web interface of the initial filtering stage.
Figure 12: Web interface of the manual correction stage.

### v2

@@
 “S-Obj” indicates single-object while “M-Obj” indicates multi-object.
-Benchmark Size Reference Images Real World Scenario Game World Scenario Attribute Temporal Pose Spatial Motion Mechanic Medium Logical Long-planing Strategic Spatial (S-Obj) (S-Obj) (S-Obj) (M-Obj) (M-Obj) (M-Obj) (M-Obj) SmartEdit [11] 219 219 RISE [52] 360 70 KRIS [41] 1,267 50 UniREditBench 2,700 2,700
+Benchmark Size Reference Images Real World Scenario Game World Scenario Attribute Temporal Pose Spatial Motion Mechanic Medium Logical Long-planing Strategic Spatial (S-Obj) (S-Obj) (S-Obj) (M-Obj) (M-Obj) (M-Obj) (M-Obj) SmartEdit [huang2024smartedit] 219 219 RISE [zhao2025envisioning] 360 70 KRIS [wu2025kris] 1,267 50 UniREditBench 2,700 2,700
 Recent advances in multimodal generative models have led to remarkable improvements in instruction-conditioned image editing.
-Generative models [40, 47, 50, 3, 33, 38, 44], including Step1X-Edit [20], FLUX-Kontext [2], Bagel [6], Nano Banana [7], and GPT-4o [12], have demonstrated a powerful ability to understand diverse textual instructions and generate semantically consistent image edits.
-In parallel, reinforcement learning-based training strategies [43, 18, 36, 31] are continuously advancing, further enhancing the capabilities of image editing models.
+Generative models [wu2025omnigen2, yu2025anyedit, zhang2023magicbrush, brooks2023instructpix2pix, wang2024primecomposer, wang2025dreamtext, xin2025lumina], including Step1X-Edit [liu2025step1x], FLUX-Kontext [flux_kontext], Bagel [deng2025bagel], Nano Banana [Gemini-2.5-Flash-Image], and GPT-4o [hurst2024gpt], have demonstrated a powerful ability to understand diverse textual instructions and generate semantically consistent image edits.
+In parallel, reinforcement learning-based training strategies [xiao2025mindomni, li2025uniworld, wang2025pref, wang2024lift] are continuously advancing, further enhancing the capabilities of image editing models.
 With these rapid developments, the need for a more comprehensive benchmark to evaluate model editing capabilities across different aspects has become increasingly essential.
-Early benchmarks [20, 46] focus on local details or global stylistic changes, e.g., style transfer, color alteration, and object removal.
-However, they fail to cover editing tasks that require models to perform implicit reasoning [49, 8], which are commonly used in real-life applications.
+Early benchmarks [liu2025step1x, ye2025imgedit] focus on local details or global stylistic changes, e.g., style transfer, color alteration, and object removal.
+However, they fail to cover editing tasks that require models to perform implicit reasoning [zhang2025r, he2025reasoning], which are commonly used in real-life applications.
 As illustrated in Fig.
 1, when editing instructions involving real-world or human-defined game rules, current models often generate results that lack physical plausibility.
-To this end, recent efforts have introduced reasoning-aware evaluation across temporal, spatial, and logical dimensions [52], and proposed a knowledge-grounded taxonomy assessing factual, conceptual, and procedural knowledge types [41].
+To this end, recent efforts have introduced reasoning-aware evaluation across temporal, spatial, and logical dimensions [zhao2025envisioning], and proposed a knowledge-grounded taxonomy assessing factual, conceptual, and procedural knowledge types [wu2025kris].
 Despite their effectiveness, these benchmarks still face two significant challenges: (1) they primarily focus on single-object attribute changes in realistic scenarios, neglecting multi-object interactions and game-world scenarios involving human-defined rules (see Tab.
 1).
-This narrow scope restricts their ability to evaluate how effectively models generalize across a wider range of complex reasoning contexts; Additionally, (2) they mainly rely on textual reference to evaluate the generated images [52, 41], which may lead to systematic misjudgments, especially in complex reasoning-based editing scenarios (see Fig.
+This narrow scope restricts their ability to evaluate how effectively models generalize across a wider range of complex reasoning contexts; Additionally, (2) they mainly rely on textual reference to evaluate the generated images [zhao2025envisioning, wu2025kris], which may lead to systematic misjudgments, especially in complex reasoning-based editing scenarios (see Fig.
 2).
@@
 2, in contrast to existing work that relies solely on textual references for evaluation, we introduce additional reference GT images to facilitate direct visual comparison with the generated image.
-By utilizing the visual cues provided by the reference image, the evaluator is able to able to more accurately and reliably assess the alignment of the generated image with the given instruction, as shown in Fig.
+By utilizing the visual cues provided by the reference image, the evaluator is able to more accurately and reliably assess the alignment of the generated image with the given instruction, as shown in Fig.
 2 (b).
@@
 Finally, all resulted textual descriptions are directly used to generate pairs of original and edited image.
-(b) For game-world scenarios, we first design diverse game problems, and then use Python programs to generate image pairs, instructions, and textual reference of edited effects, ensuring both logical and visual correctness in these rule-intensive scenarios [16, 27].
+(b) For game-world scenarios, we first design diverse game problems, and then use Python programs to generate image pairs, instructions, and textual reference of edited effects, ensuring both logical and visual correctness in these rule-intensive scenarios [li2024vcbench, tong2025code2logic].
 Ultimately, all data samples in UniREditBench undergo VLM-based filtering and human inspection to ensure their reliability and accuracy.
@@
 3.
-To validate its reliability and effectiveness, we fine-tune the Bagel [6] on this dataset, resulting in UniREdit-Bagel.
-Experimental results demonstrate that the fine-tuned model achieves substantial improvements on both UniREditBench and other out-of-distribution benchmarks [52, 41].
+To validate its reliability and effectiveness, we fine-tune the Bagel [deng2025bagel] on this dataset, resulting in UniREdit-Bagel.
+Experimental results demonstrate that the fine-tuned model achieves substantial improvements on both UniREditBench and other out-of-distribution benchmarks [zhao2025envisioning, wu2025kris].
 Additionally, through comprehensive evaluation of both open- and closed-source editing models on our UniREditBench, we reveal their strengths and weaknesses across diverse reasoning-based scenarios.
@@
 Instruction-based image editing models aim to bridge semantic understanding of instructions with accurate visual manipulation.
-Traditional methods perform editing by altering the diffusion trajectory without requiring additional training, including partial denoising from intermediate SDE steps [21], cross-attention control [10, 30], mask-guided blending [38, 1, 33], CLIP- or diffusion-guided manipulation [14], and latent inversion for fidelity preservation [32, 13].
-Besides, several studies employ visual-language models (VLMs) to provide prompts, spatial priors, or synthetic supervision to guide a generative editing model [3, 50, 51, 9].
-Recent unified frameworks aim to use a single model for both image understanding and editing in a complementary direction [44, 40, 29].
-For instance, Bagel [6] features a think mode that produces reasoning text prior to editing to enhance instruction fidelity and consistency.
+Traditional methods perform editing by altering the diffusion trajectory without requiring additional training, including partial denoising from intermediate SDE steps [meng2021sdedit], cross-attention control [hertz2022prompt, wang2023enhancing], mask-guided blending [wang2025dreamtext, avrahami2022blended, wang2024primecomposer], CLIP- or diffusion-guided manipulation [kim2022diffusionclip], and latent inversion for fidelity preservation [wang2024magicface, kawar2023imagic].
+Besides, several studies employ visual-language models (VLMs) to provide prompts, spatial priors, or synthetic supervision to guide a generative editing model [brooks2023instructpix2pix, zhang2023magicbrush, zhang2024hive, he2024freeedit, han2025controlthinker].
+Recent unified frameworks aim to use a single model for both image understanding and editing in a complementary direction [xin2025lumina, wu2025omnigen2, wang2024emu3].
+For instance, Bagel [deng2025bagel] features a think mode that produces reasoning text prior to editing to enhance instruction fidelity and consistency.
 While effective, current methods still face challenges with complex reasoning-based editing, underscoring the need for comprehensive benchmarks to assess their performance across various reasoning scenarios.
 Reasoning-based Benchmarks for Image Generation and Editing.
-In T2I generation, several benchmarks [34, 25, 17, 22] have been developed to assess the reasoning capabilities of models in generating images.
-For example, WISE [22] focuses on assessing models’ world knowledge, such as cultural and physical understanding, while UniGenBench++ [34] unifies semantic generation evaluation, covering 10 primary dimensions and 27 sub-dimensions, such as logic reasoning, relational understanding, supporting multilingual and varying-length assessments.
-In image editing evaluation, recent reasoning-based benchmarks like RISEBench [52] aim to examine temporal, spatial, and logical editing capabilities of editing models.
-Besides, KRISBench [41] introduces a knowledge-grounded taxonomy covering factual, conceptual, and procedural types.
+In T2I generation, several benchmarks [wang2025unigenbench++, sun2025t2i, li2025easier, niu2025wise] have been developed to assess the reasoning capabilities of models in generating images.
+For example, WISE [niu2025wise] focuses on assessing models’ world knowledge, such as cultural and physical understanding, while UniGenBench++ [wang2025unigenbench++] unifies semantic generation evaluation, covering 10 primary dimensions and 27 sub-dimensions, such as logic reasoning, relational understanding, supporting multilingual and varying-length assessments.
+In image editing evaluation, recent reasoning-based benchmarks like RISEBench [zhao2025envisioning] aim to examine temporal, spatial, and logical editing capabilities of editing models.
+Besides, KRISBench [wu2025kris] introduces a knowledge-grounded taxonomy covering factual, conceptual, and procedural types.
 However, these benchmarks primarily focus on single-object knowledge and attribute reasoning.
 We suppose that extending evaluation to multi-object interactions and scenarios governed by human-defined rules is a crucial next step.
-As for image quality evaluation [45, 15], recent works like UnifiedReward [37, 35] adopt the “VLM-as-a-judge” paradigm, leveraging the powerful capabilities of VLMs to score and provide explanatory judgments.
+As for image quality evaluation [xu2023imagereward, kirstain2023pick], recent works like UnifiedReward [wang2025unified, wang2025unified-think] adopt the “VLM-as-a-judge” paradigm, leveraging the powerful capabilities of VLMs to score and provide explanatory judgments.
 In image editing tasks, evaluation is more challenging because the evaluator needs to assess not only image quality but also understand complex editing instructions and final edited effects.
-Most studies like RISEBench and KRISBench utilize the property model [12], to rate instruction following, temporal consistency, and image quality.
+Most studies like RISEBench and KRISBench utilize the property model [hurst2024gpt], to rate instruction following, temporal consistency, and image quality.
 Despite effectiveness, their evaluation relies solely on textual references, which may lead to systematic misjudgments in complex reasoning tasks.
@@
 Specifically, for each sample, this pipeline evaluates three core dimensions:
-  • Instruction Following measures how accurately the generated image reflects the input instruction, focusing on whether the explicit effect of the edits is properly manifested.
+  • Instruction Following measures how accurately the generated image reflects the input instruction, focusing on whether explicit effect of the edit is properly manifested.
 Here, the VLM compares the output image $G$ against both the textual reference of edited effect $R_{t}$ and the corresponding reference GT image $R_{i}$ to verify compliance:
@@
     $S_{\mathrm{VC}}=\mathrm{VLM}(O,I,G)$
-  • Visual Quality evaluates the realism and perceptual integrity of the generated output, checking for artifacts, distortions, and physical or logical implausibility in the final image:
+  • Visual Quality evaluates the realism of the generated output, checking for artifacts, distortions, and physical or logical implausibility in the final image:
     $S_{\mathrm{VQ}}=\mathrm{VLM}(G)$
-    We choose GPT-4.1 [12] as VLM evaluator.
-Each score $S$ ranges from 1 to 5, following prior detailed scoring guidelines [52, 41].
+    We choose GPT-4.1 [hurst2024gpt] as VLM evaluator.
+Each score $S$ ranges from 1 to 5, following prior detailed scoring guidelines [zhao2025envisioning, wu2025kris].
 Finally, the overall evaluation score aggregates these via weighted sum:
@@
 […]
-FLUX-Kontext-Pro 45.68 45.65 42.85 40.17 43.59 43.54 44.03 49.80 40.07 44.36 43.72
-Seedream4.0 69.54 73.13 67.88 62.40 68.24 39.27 43.54 43.79 51.91 44.63 57.05
-Wan2.5 74.13 69.92 64.23 65.16 68.36 63.39 46.13 56.32 52.24 54.52 60.59
-Nano Banana 77.54 78.45 72.41 77.39 76.45 66.26 56.83 56.35 65.85 61.32 68.38
-GPT-4o 83.27 83.93 80.43 77.62 81.31 78.82 50.28 66.02 65.66 65.19 71.64
+FLUX-Kontext-Pro 47.35 47.16 44.37 41.44 45.00 49.12 48.58 51.16 40.49 46.52 45.77
+Seedream4.0 67.98 72.93 65.07 59.49 66.22 38.87 42.16 44.09 51.65 45.38 55.77
+Wan2.5 74.66 69.74 63.51 62.87 67.23 63.73 47.13 55.00 54.93 52.67 61.36
+Nano Banana 77.10 78.88 71.86 74.70 75.22 66.74 56.11 56.83 64.91 60.39 68.26
+GPT-4o 82.67 84.75 79.81 77.40 81.01 77.73 51.44 67.61 63.79 62.07 73.39
 […]
-MagicBrush 43.10 46.24 43.88 47.70 45.23 63.43 33.77 30.83 35.17 40.80 40.98
-Omnigen2 54.18 57.75 52.87 53.52 54.58 70.78 27.61 37.11 24.69 40.05 43.96
-Step1X-Edit 59.50 57.37 56.07 56.75 57.42 62.90 34.17 44.63 44.03 46.43 50.12
-Bagel-Think 61.27 59.83 53.95 57.11 58.04 65.65 43.46 43.82 40.18 48.28 51.25
-Qwen-Image-Edit 75.19 73.16 71.80 67.54 71.92 57.03 34.80 47.37 37.70 44.22 56.46
-UniREdit-Bagel (Ours) 79.24 78.61 76.43 71.81 76.52 84.93 73.98 86.14 84.01 82.27 78.87
+MagicBrush 43.97 46.09 42.93 46.64 44.69 63.58 33.59 30.72 35.31 36.85 40.77
+Omnigen2 55.47 58.27 51.44 50.70 53.69 70.28 27.50 36.25 24.31 33.14 43.41
+Lumina-DiMOO 52.84 52.31 50.31 50.83 51.44 61.23 36.96 39.57 53.09 45.61 48.54
+Step1X-Edit 59.69 56.36 53.85 54.84 55.93 65.68 34.89 47.01 43.89 44.00 50.15
+Bagel-Think 61.45 59.51 52.65 55.68 56.80 66.29 43.23 43.00 41.30 45.10 50.96
+DreamOmni2 63.52 58.35 52.68 54.02 56.64 72.42 42.17 48.80 48.09 48.98 52.81
+UniWorld-V2 72.96 69.37 63.65 61.69 66.55 49.27 40.00 53.41 37.53 43.19 54.87
+Qwen-Image-Edit 75.68 73.03 70.59 64.67 70.95 56.73 36.63 48.80 37.68 41.92 56.52
+UniREdit-Bagel (Ours) 76.73 77.80 76.57 71.44 75.74 84.90 72.83 84.88 83.72 80.48 78.15
 3.4 Multi-Scenario Data Synthesis
@@
 Specifically, (1) this process begins with hand-crafted textual triples that describe the original image, the editing instruction, and the textual reference of edited effect (a reasoning-based narrative of the anticipated outcome).
-Next, we use the powerful VLM [5], to expand this initial set into a large corpus of text triples.
-Subsequently, (2) these curated textual triples are input to GPT-4o [12] to synthesize the original and edited images in alignment with the described textual reference of edited effect.
-(3) Finally, in the quality filtering stage, VLM [5] is used to assess the generated images based on visual fidelity, instruction alignment, and potential hallucination risks.
+Next, we use the powerful VLM [comanici2025gemini], to expand this initial set into a large corpus of text triples.
+Subsequently, (2) these curated textual triples are input to GPT-4o [hurst2024gpt] to synthesize the original and edited images in alignment with the described textual reference of edited effect.
+(3) Finally, in the quality filtering stage, VLM [comanici2025gemini] is used to assess the generated images based on visual fidelity, instruction alignment, and potential hallucination risks.
 Additionally, it generates reasoning chain-of-thought (CoT) text for each qualified instance, ensuring the production of high-quality, reasoning-based image editing training data.
 In Game-World Scenario, game states are inherently well-suited to be represented as structured reasoning-based editing data, where instructions can naturally be solved using Python code.
-Inspired by Game-RL [48], (1) we first design a diverse collection of game-based problems and develop corresponding Python programs tailored to each category.
+Inspired by Game-RL [tong2025code2logic], (1) we first design a diverse collection of game-based problems and develop corresponding Python programs tailored to each category.
 (2) Then, these programs automatically generate paired original and edited images, along with instructions, textual reference effects, and programmatic CoT reasoning traces.
@@
 Finally, quality filtering are applied to ensure the integrity and reliability of the data.
-Overall, this multi-scenario data synthesis pipeline generates our UniREditBench, a unified reasoning-based image editing benchmark, and UniREdit-Data-100K, a large-scale synthetic dataset with high-quality CoT annotations.
-We will provide a detailed elaboration of this dataset in the following section.
+Overall, this multi-scenario data synthesis pipeline generates UniREditBench, a unified reasoning-based image editing benchmark, and UniREdit-Data-100K, a large-scale dataset with high-quality CoT annotations.
+We detail the elaboration of this dataset in the next section.
 4 UniREdit-Data-100K
@@
 UniREdit-Data-100K is designed with an emphasis on balance and diversity, ensuring that each reasoning category contains over 4,000 instances to effectively support model training across a wide range of editing tasks.
-It is divided into two primary scenario types: (i) Real-World Scenario, which captures natural object attributes and complex multi-object interactions, and (ii) Game-World Scenario, presenting structured, rule-based editing challenges, such as puzzles and strategic planning games.
+It is divided into two primary scenarios: (i) Real-World Scenario, which captures natural object attributes and complex multi-object interactions, and (ii) Game-World Scenario, presenting structured, rule-based editing challenges, such as puzzles and strategic planning games.
 We visualize the word cloud for both real-world and game-world subsets in Fig.
@@
 5 (b).
-These visualizations highlight the extensive vocabulary of our dataset that captures the diverse visual attributes, as well as its broad coverage across various categories.
+These visualizations highlight the extensive vocabulary as well as the broad coverage across various categories of our dataset.
 4.2 UniREdit-Bagel
-To further validate the effectiveness of our dataset, we use it to fine-tune Bagel [6], a unified understanding and generative model.
+To further validate the effectiveness of our dataset, we use it to fine-tune Bagel [deng2025bagel], a unified understanding and generative model.
 Specifically, each training sample consists of the input image $O$, an editing instruction $I$, a stepwise CoT text $C$ that grounds the edit effects step by step, and the target edited image $G$.
@@
 $\mathcal{L}_{\text{text}}=-\sum^{T}_{t}\log p_{\theta}\big(y_{t}\mid y_{<t},O,I\big).$
-For image generation, we supervise the latent flow-matching loss [19] between the VAE latents of $O$ and $G$, conditioned on $(O,I,C)$:
+For image generation, we supervise the latent flow-matching loss [lipman2022flow] between the VAE latents of $O$ and $G$, conditioned on $(O,I,C)$:
 $\mathcal{L}_{\text{img}}=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\;\big\|u_{\theta}(z_{t},t\,;\,O,I,C)-u^{\star}(z_{t},t)\big\|_{2}^{2},$
@@
 Baselines.
-We benchmark closed-source models including: GPT-4o [12], Nano Banana [7], Gemini-2.0 [5], Seedream4.0 [24], Wan 2.5 [28], and FLUX-Kontext-Pro [2], as well as open-source models including: Bagel [6], Qwen-Image-Edit [39], Step1X-Edit [20], FLUX.1-Kontext-dev [2], Emu2 [26], Omnigen2 [40], Omnigen [42], HiDream-Edit [4], MagicBrush [50], and AnyEdit [47].
+We benchmark closed-source models including: GPT-4o [hurst2024gpt], Nano Banana [Gemini-2.5-Flash-Image], Gemini-2.0 [comanici2025gemini], Seedream4.0 [seedream4], Wan 2.5 [wan2025wan], and FLUX-Kontext-Pro [flux_kontext], as well as open-source models including: Bagel [deng2025bagel], Qwen-Image-Edit [qwen-image], Step1X-Edit [liu2025step1x], FLUX.1-Kontext-dev [flux_kontext], Emu2 [sun2024generative], Omnigen2 [wu2025omnigen2], Omnigen [xiao2025omnigen], HiDream-Edit [cai2025hidream], MagicBrush [zhang2023magicbrush], and AnyEdit [yu2025anyedit].
 Training and Evaluation.
-We train all components of Bagel [6] except the VAE model for 5,000 iterations using the Adam optimizer and the cosine learning rate scheduler on UniREdit-Data-100K.
-The scheduler includes 500 warm-up steps, with the peak learning rate of 2e-5 and minimum learning rate of 1e-6.
+We train all Bagel [deng2025bagel] components, except the VAE, for 5,000 iterations on UniREdit-Data-100K using the Adam optimizer and a cosine learning-rate schedule with 500 warm-up steps, a peak learning rate of $2\times 10^{-5}$, and a minimum learning rate of $10^{-6}$.
+The loss weights are set to $\lambda_{\text{text}}=2$ and $\lambda_{\text{img}}=1$.
 During inference, we use the official inference settings provided by Bagel.
@@
 As shown in Tab.
-3.3, among closed-source models, GPT-4o achieves the highest average performance across all scenarios, with Nano Banana showing comparable capability.
+3.3, among closed-source models, GPT-4o achieves the highest average performance across all scenarios, with Nano Banana performing comparably.
 Wan2.5 delivers balanced results on real-world tasks but lags on game scenarios that require strategic reasoning.
-Besides, Seedream4.0 performs reasonably well on the structure transform dimension yet encounters challenges in game scenarios.
+Besides, Seedream4.0 is competitive on structure transform yet encounters challenges in game scenarios.
 Among open-source baselines, Qwen-Image-Edit performs strongly on real-world tasks such as attribute modification and structure transform.
@@
 In contrast, UniREdit-Bagel excels in both fulfilling the instruction and maintaining the coherence of unrelated content.
-Table 3: Out-of-distribution quantitative performance comparison on RISEBench [52]. GPT-4.1 is used as the evaluator. Best scores are in bold.
+Table 3: Out-of-distribution quantitative performance comparison on RISEBench [zhao2025envisioning]. GPT-4.1 is used as the evaluator. Best scores are in bold.
 Models Temporal Causal Spatial Logical Overall
@@
 6 Conclusion
-This paper introduces UniREditBench, a unified reasoning-based benchmark for image editing with broader evaluation dimension coverage and a reliable dual-reference evaluation pipeline.
-Additionally, we design a multi-scenario data synthesis pipeline and release UniREdit-Data-100K, a large-scale dataset with high-quality chain-of-thought (CoT) annotations.
-To demonstrate its effectiveness, we fine-tune Bagel on this dataset, resulting in UniREdit-Bagel, which achieves significant improvements both quantitatively and qualitatively.
-Through comprehensive benchmarking of both open-source and closed-source image editing models, we highlight their strengths and weaknesses across various aspects.
+This paper presents UniREditBench, a unified reasoning-based benchmark for image editing with broad dimension coverage and a robust dual-reference evaluation protocol.
+We further introduce a multi-scenario synthesis pipeline and release UniREdit-Data-100K, a large-scale dataset with high-quality CoT annotations.
+To demonstrate its effectiveness, we fine-tune Bagel on this dataset and yields UniREdit-Bagel, which achieves substantial quantitative and qualitative gains.
+Comprehensive benchmarking of open-source and closed-source image editing models reveals their strengths and weaknesses across various aspects.
 Supplementary Material
@@
 These duplicate samples are then filtered out to ensure data diversity.
-  • Semantic-Similarity Deduplication: We use a sentence-transformers [23] model to extract sentence embeddings for both the original image description and edit instruction.
+  • Semantic-Similarity Deduplication: We use a sentence-transformers model to extract sentence embeddings for both the original image description and edit instruction.
 We then compute the pairwise similarity between these embeddings.
@@
 Figure 7: Benchmarking result visualization. (a) Closed-source model comparison; (b) Open-source model comparison.
-Table 4: Quantitative comparisons on KRISBench [41]. GPT-4.1 is used as the evaluator. Best scores are in bold while second-best is underlined.
+Table 4: Quantitative comparisons on KRISBench. GPT-4.1 is used as the evaluator. Best scores are in bold while second-best is underlined.
 Model Attribute Spatial Temporal Social Natural Logical Instruction Overall
@@
 […]
-FLUX-Kontext-Pro 34.75 56.62 54.95 36.36 42.15 42.52 43.89 38.01 42.33 43.54 47.45 34.09 50.55 34.88 58.44 26.88 46.70 52.90 43.72
-Seedream4.0 64.77 74.32 80.19 66.07 59.48 64.68 79.48 61.58 63.21 39.27 46.28 43.00 41.34 48.55 38.43 68.75 54.15 33.43 57.05
-Wan2.5 72.97 75.30 79.97 59.86 64.51 66.56 61.60 63.67 66.66 63.39 44.67 53.73 40.00 58.54 60.59 45.94 65.17 47.47 60.59
-Nano Banana 75.37 79.72 85.55 71.35 70.65 73.36 73.22 79.44 75.34 66.26 61.43 54.65 54.40 64.65 40.92 91.99 62.65 50.05 68.38
-GPT-4o 83.55 82.98 92.16 75.70 76.10 76.32 88.88 78.97 76.28 78.82 58.88 44.95 47.02 64.55 63.80 68.62 82.27 49.77 71.64
+FLUX-Kontext-Pro 37.55 57.15 56.79 37.52 41.46 44.62 47.03 40.56 42.32 49.12 53.18 38.12 54.43 32.45 61.23 27.80 48.45 53.87 45.77
+Seedream4.0 64.18 71.78 79.98 65.87 56.10 61.60 77.52 59.15 59.83 38.87 45.47 40.87 40.13 48.05 38.17 68.72 55.27 32.90 55.77
+Wan2.5 73.78 75.54 79.95 59.53 62.67 64.38 63.49 60.83 64.91 63.73 43.44 56.57 41.38 58.38 58.57 47.86 64.03 45.97 61.36
+Nano Banana 75.82 78.39 86.12 71.63 71.37 71.07 73.14 77.58 71.82 66.74 59.97 54.33 54.04 64.13 39.80 90.81 62.50 51.15 68.26
+GPT-4o 83.83 81.52 92.18 77.33 75.86 74.98 88.60 78.98 75.81 77.73 58.96 45.60 49.75 64.62 58.97 67.78 83.73 51.49 73.39
 […]
-MagicBrush 35.65 50.55 48.34 44.13 45.07 47.83 38.73 47.88 47.52 63.43 44.74 26.26 30.30 34.12 31.05 40.33 32.65 29.00 40.98
-Omnigen2 50.32 58.03 67.73 47.77 49.90 57.63 51.09 55.05 51.98 70.78 51.17 1.50 30.17 23.07 46.17 4.85 39.63 34.58 43.96
-Step-1X-Edit 51.82 67.18 63.29 51.45 54.98 62.18 51.03 55.72 57.78 62.90 33.62 35.40 33.48 43.02 49.92 39.17 54.53 34.73 50.12
-Bagel-Think 58.38 64.15 63.38 56.27 54.40 58.63 48.82 56.47 57.75 65.65 47.30 42.12 40.97 47.80 40.35 32.40 48.83 38.80 51.25
-Qwen-Image-Edit 72.08 78.31 81.75 64.57 68.78 69.51 77.12 67.83 67.25 57.03 36.58 37.79 30.02 48.47 33.00 31.63 60.02 34.71 56.46
-UniREdit-Bagel (Ours) 84.43 74.06 85.13 72.08 74.31 71.95 83.05 71.17 72.45 84.93 86.93 61.25 73.76 87.48 70.65 93.90 97.73 74.55 78.87
+MagicBrush 36.74 51.20 47.75 44.43 44.05 47.06 37.67 45.87 47.42 63.58 45.33 25.60 29.83 33.92 31.45 40.55 33.20 28.23 40.77
+Omnigen2 51.25 59.68 68.25 48.28 48.37 54.48 51.48 50.77 50.64 70.28 51.33 1.38 29.80 22.87 44.97 5.10 39.57 32.93 43.41
+Lumina-DiMOO 54.22 51.47 54.40 50.23 48.67 52.62 49.65 51.63 50.03 61.23 48.23 36.29 26.35 41.73 65.07 52.47 43.57 35.57 48.54
+Step-1X-Edit 52.57 66.82 62.15 50.58 53.07 58.68 49.80 54.23 55.45 65.68 35.05 36.38 33.25 45.49 46.97 39.20 55.94 38.08 50.15
+Bagel-Think 59.05 63.84 63.52 55.49 54.05 55.62 48.27 55.03 56.33 66.29 45.48 42.89 41.33 48.42 42.20 33.27 48.62 37.38 50.96
+DreamOmni2 61.22 65.82 67.07 49.63 49.58 54.83 53.62 54.11 53.92 72.42 38.78 42.33 45.40 45.86 46.98 51.42 56.55 41.05 52.81
+UniWorld-V2 71.45 74.47 83.19 55.55 58.53 65.47 66.95 62.20 61.17 49.27 34.44 42.90 42.65 49.67 34.73 28.20 71.59 35.22 54.87
+Qwen-Image-Edit 74.60 76.76 81.45 64.62 66.48 68.78 76.50 64.65 64.69 56.73 38.42 41.02 30.45 48.52 33.60 30.93 61.38 36.21 56.52
+UniREdit-Bagel (Ours) 81.08 72.39 84.78 70.81 73.85 72.73 83.13 71.27 71.62 84.90 87.77 57.68 73.05 86.80 71.16 93.20 98.30 71.47 78.15
 A.3 Human Inspection
@@
   • Manual Correction: Annotators make refinements to the textual reference effect that are only slightly incorrect, ensuring alignment and accuracy.
-Two web interfaces are shown in Figs. 11 and  12.
+Two web interfaces are shown in Figs. 12 and  13.
 Appendix B Detailed Benchmarking Results
 We provide detailed benchmarking results on our UniREditBench for each category in Tab. A.2.
+Figure 8: Overall performance on UniREditBench of models trained on different size of UniREdit-Data.
 Appendix C More Quantitative Results
@@
 Appendix D More Qualitative Comparison Results
-We provide additional qualitative comparisons on UniREditBench in Fig. 8 and 9, and comparisons on RISEBench in Fig. 10.
+We provide additional qualitative comparisons on UniREditBench in Fig. 9 and 10, and comparisons on RISEBench in Fig. 11.
+Appendix E Effect of Training Set Size on Overall Score
+We provide overall performance comparison on UniREditBench of models trained on difference size of UniREdit-Data, which is visualized in Fig. 8.
 Figure 9: Qualitative editing result comparison on UniREditBench.
@@
 Our UniREdit-Bagel demonstrates significant superiority in both instruction following and visual quality compared with state-of-the-art closed-sourced and open-sourced models.
-Appendix E Ethical statement
+Appendix F Ethical statement
 In this work, we affirm our commitment to ethical research practices and responsible innovation.